# Ultra-Slow-Roll Inflation on the Lattice: Backreaction and Nonlinear Effects

### v2

@@ Abstract
Violating the slow-roll regime during the final stages of inflation can significantly enhance curvature perturbations, a scenario often invoked in models producing primordial black holes and small-scale scalar induced gravitational waves. When perturbations are enhanced, one approaches the regime in which tree-level computations are insufficient, and nonlinear corrections may become relevant. In this work, we conduct lattice simulations of ultra-slow-roll (USR) dynamics to investigate the significance of nonlinear effects, both in terms of backreaction on the background and in the evolution of perturbations. Our systematic study of various USR potentials reveals that nonlinear corrections are significant when the tree-level curvature power spectrum peaks at $\mathcal{P}^{\rm max}_{\zeta}={\cal O}(10^{-3})-{\cal O}(10^{-2})$, with 5%--10% corrections. Larger enhancements yield even greater differences. We find a simple universal relation between simulation and tree-level quantities $\dot{\phi}=\dot{\phi}_{\rm tree}\left(1+\sqrt{\mathcal{P}^{\rm max}_{\zeta,\rm tree}}\right)$ at the end of the USR phase, which is valid in all cases we consider. Additionally, we explore how nonlinear interactions during the USR phase affect the clustering and non-Gaussianity of scalar fluctuations, crucial for understanding the phenomenological consequences of USR, such as scalar-induced gravitational waves and primordial black holes. Our findings demonstrate the necessity of going beyond leading order perturbation theory results, through higher-order or non-perturbative computations, to make robust predictions for inflation models exhibiting a USR phase.
@@ I Introduction
However, in certain models, the inflaton potential features a flat region or shallow minimum where the deceleration becomes dominant, and the field’s velocity can decrease exponentially, in what is called an “ultra-slow-roll” (USR) phase Ivanov et al. (1994); Kinney (1997); Inoue and Yokoyama (2002); Kinney (2005); Martin et al. (2013); Motohashi and Hu (2017). Perturbations generated during USR are far from scale-invariant, displaying a large amplification. These perturbations may lead to interesting phenomenological consequences, such as the emission of scalar-induced gravitational waves (SIGW) Tomita (1975); Matarrese et al. (1994); Acquaviva et al. (2003); Mollerach et al. (2004); Ananda et al. (2007); Baumann et al. (2007); Domènech (2021), and/or seed primordial black hole (PBH) formation Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974); Carr (1975); Chapline (1975). PBHs are a longstanding candidate for dark matter, potentially accounting for all dark matter or acting as seeds for supermassive black holes, see Refs. Carr et al. (2021); Green and Kavanagh (2021) for recent reviews. The blooming field of GW astronomy will provide unprecedented observations of small-scale GWs, being able to constrain the early universe (see Achúcarro et al. (2022); Auclair et al. (2023); Bagui et al. (2023) for recent reviews).
A very relevant question is whether the dynamics of USR are consistent with perturbativity (see Ref. Kristiano and Yokoyama (2024a) for a recent review). From a technical standpoint, the dimensionless power spectrum of curvature perturbations, $\mathcal{P}_{\zeta}(k)$, is usually calculated within linear, i.e. tree-level, perturbation theory. However, as linear perturbations can become large (often reaching $\sqrt{\mathcal{P}_{\zeta}(k)}\simeq 0.1$), curvature perturbations are expected to be sensitive to nonlinear interactions. These nonlinear effects have various consequences. First, they introduce non-zero higher-order cumulants in addition to the variance, with important implications for the PBH formation, see e.g. Franciolini et al. (2018); Taoso and Urbano (2021). Second, through loops, they modify the variance itself compared to the result obtained from the free theory. Lastly, they lead to a backreaction on the background dynamics. In this paper we investigate the relevance of these effects, considering motivated scenarios built with the reverse engineering approach proposed in Ref. Franciolini and Urbano (2022), by performing numerical lattice simulations able to capture the dynamics of the inflaton non-perturbatively.
@@ II.1 Inflationary background
The inflationary background can be described by modelling the evolution of the Hubble rate $H\equiv\dot{a}/a$, where $a$ is the scale factor during inflation. This is dictated by dynamical equations relating $H$ to the Hubble parameters, which are
@@
where $\dot{H}=dH/dt$ is the cosmic-time derivative of $H$ while $N$, defined such as $dN=Hdt$, is the number of $e$-folds.111 Notice that a different definition of the Hubble parameter $\eta$ is sometimes used, and defined as $\epsilon_{2}\equiv\dot{\epsilon}/H\epsilon=-2(\eta-\epsilon)$.
@@
where the parameter $\delta N$ controls the width of the two transitions at $N_{\rm in}$ and $N_{\rm end}$. We assume that the second Hubble parameter $\eta$ is negligible in the initial slow-roll phase $\eta_{\rm I}\simeq 0$, consistently with the CMB requirement in the earlier stage of inflation. Once the parameters in Eq. (2) are fixed, one can solve for the evolution of $\epsilon$ and $H$ as a function of number of $e$-folds using (1). In all cases, we arbitrarily set $N_{\rm in}=0$. In Fig. 1, we show the evolution of $\epsilon$ and $\eta$ for the scenarios considered in this work, which can be divided into three cases, depending on the value of $\eta$ during USR: i) the approximate Wands duality (case I); ii) the repulsive (case II); iii) the attractive (case III). The choice of these names will become clear in the following.
@@
Figure 1: Evolution of $\eta(N)$ (top panel) and $\epsilon(N)$ (bottom panel) as a function of number of $e$-folds considered in this work. Solid, dashed and dotted line indicate different cases, corresponding to the approximate Wands duality (WD), repulsive and attractive case, respectively. See the main text for more details. Different colors (from purple to yellow) correspond to different USR durations $N_{\rm end}-N_{\rm in}$, leading to a tree-level curvature power spectrum at its maximum ranging from ${\cal P}_{\zeta}^{\rm max}=10^{-4}$ to ${\cal P}^{\rm max}_{\zeta}=1$. In all cases, we fixed $N_{\rm in}=0$.
@@
In Tab. 1 we report parameters we adopt in the various cases. We arbitrarily fix $\delta N=0.5$, in order to have relatively smooth transitions in and out of the USR phase. Additionally, we choose values of $\eta_{\rm II}$ that are in all cases larger than $\simeq 3/2$, as required to enter in the USR regime. Finally, we set $\eta_{\rm III}=-0.5$, as a negative value is required in order to end inflation by raising $\epsilon$ again to reach ${\cal O}(1)$. [^2] For simplicity, we use the terminology “slow-roll” for the third phase, by contrast with the preceding USR phase, despite the relatively large value of $\eta_{\rm III}$. In case I, we approximately realise a scenario that respects the Wands duality (WD) Wands (1999), stating that the same tree-level spectrum of perturbations is obtained during phases characterised by a constant $\eta$ or $3-\eta$. In case I, this property is respected between phase II and III, and therefore it leads to a tree-level curvature power spectrum featuring a single smooth peak that can be fitted with a broken power-law. This happens because the spectrum of modes exiting the Hubble scale during USR follows $k^{2(3-\eta_{\rm II})}\sim 1/k$, while $k^{2\eta_{\rm III}}\sim 1/k$ in the subsequent phase Karam et al. (2023); Franciolini and Urbano (2022). This property can also be derived by inspecting the effective mass of scalar perturbations when $\epsilon$ is negligible. We will come back to this point in Sec. II.3.2. As we will see, the nearly constant effective mass term will also be accompanied by negligible higher-order derivatives of the potential, rendering the theory nearly free from self-interactions at these times. In the cases II and III, $\eta_{\rm II}$ is taken to be smaller or larger than the WD value, leading to different phenomenology for the scalar perturbations.
@@ II.2 Inflaton potential
where in the second equation we consider the minus sign having in mind a large-field model in which the field value decreases as inflation proceeds. Combining $V(N)$ and $\phi(N)$, we reconstruct the profile $V(\phi)$ of the inflationary potential in field space. In all cases, we arbitrarily choose $\phi(N_{\rm CMB})=3.5$ and, by fixing the amplitude of the power spectrum at large scales to be compatible with CMB data, i.e. ${\cal P}_{\zeta}=2.1\cdot 10^{-9}$, we find $V(N_{\rm CMB})=3.2\cdot 10^{-9}$. The initial condition satisfying the SR attractor at $N_{\rm CMB}$ require $\phi(N_{\rm CMB})=3.5$ and ${\rm d}\phi(N_{\rm CMB})/{\rm d}N=-0.11$. Eq. (3a) shows the convenience of modelling the inflationary dynamics directly at the level of $\eta$ instead of $V(\phi)$. This is because the Hubble parameters enters at the exponent of the definition of $V(N)$, and thus allow for a much finer control on power spectral features when performing the reverse engineering procedure.
In Fig. 2, we show the reconstructed potential for the various scenarios considered in this work, zooming in the shallower region of the potential that controls the USR dynamics. We also show the first and second derivatives of the inflaton potential, as they will be relevant for the subsequent discussion. It is interesting to notice that, in case II and III, sharp changes of the second derivative of the potential are obtained near the end of the feature, corresponding to the expected end of the USR phase. This will lead to large higher-order derivatives of the potential, crucially controlling the coupling of the theory. Notice that the potentials of case II are monotonic. As for the ones of case I and III, they present a local minimum followed by a local maximum (in the direction of decreasing $\phi$). However, the minimum is not deep enough to lead to a trapping of the inflaton Atal et al. (2019, 2020); Inomata et al. (2022); Escrivà et al. (2023); Caravano et al. (2024) (contrary to what happens in other contexts, see e.g. Caravano et al. (2024) in a model with oscillations in the potential Inomata et al. (2023)), with the exception of the most extreme situation of case III with $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=1$, as we will discuss below.
@@ II.3 Tree-level curvature power spectrum
Once the background evolution is defined, we can calculate the spectrum of gauge-invariant comoving curvature perturbations generated during inflation. As long as the slow-roll approximation holds, this spectrum is given by ${\cal P}_{\zeta}(k)={H^{2}}/{8\pi^{2}\epsilon}$ where the Hubble parameters are evaluated at the Hubble crossing of mode $k$. This hints towards the general expectation of exponentially enhanced perturbations during the USR phase, where $\epsilon$ decays exponentially.
@@
We solve the SM equation with sub-Hubble Bunch-Davies initial conditions for modes deep within the Hubble sphere, i.e. when $N$ is smaller than the Hubble crossing epoch $N_{k}$ s.t. $k=a(N_{k})H(N_{k})$. This is implemented on the SM variable as
@@
To illustrate the typical dynamics, Fig. 3 shows the evolution of curvature perturbations as a function of the number of $e$-folds for Case I with $\Delta N=2$ (corresponding to a very good approximation with $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-2}$). In the left panel, we show modes crossing the Hubble sphere before, during and right after the USR phase. While each mode experiences a growth due to the decaying mode changing behavior during USR, they all freeze out soon after USR ends and the subsequent SR evolution begins. The lattice simulations track the evolution of the system until the relevant modes have frozen. In the right panel, we indicate where these modes sits in the curvature power spectrum at the end of inflation. We normalize the x-axis with respect to $k_{\rm ref}$, which is the mode crossing the Hubble scale at $N_{\rm in}$, i.e. $k_{\rm ref}=a(N_{\rm in})H(N_{\rm in})$.
@@ II.3.1 On the irrelevance of gravitational backreaction
We can inspect the relevance of the gravitational backreaction by focusing on the SM equation (6). By neglecting gravitational backreaction, i.e. neglecting the lapse and shift perturbations in the spatially flat gauge, one obtains a different mass term, namely Eq. (7) subtracting $4\epsilon\eta-6\epsilon-2\epsilon^{2}$. In the situations of interest we have $\epsilon\ll 1$, which shows that gravitational backreaction can be safely neglected as long as $|3-\eta|\gg\epsilon$. This simply corresponds to the decoupling limit (see e.g. Creminelli et al. (2024)) in the EFT of inflation Cheung et al. (2008). This condition is satisfied in our setups and hence our lattice simulations, which do not take into account metric fluctuations, should agree with tree-level results if standard perturbation theory holds. [^3] Note that, if necessary, metric perturbation can be perturbatively taken into account in lattice simulations, as recently done in Caravano and Peloso (2024). We have confirmed this explicitely by checking the relative difference between the effective mass terms computed in the two cases. Due to the exponential decay of $\epsilon$ during the USR phase $\epsilon\sim\exp\left(-2\eta_{\rm II}N\right)$, the backreaction induced terms quickly become subdominant, while $m_{\rm eff}^{2}/H^{2}$ remain ${\cal O}(1)$.

### v3

# Ultra-Slow-Roll Inflation on the Lattice I: Backreaction and Nonlinear Effects

@@ Abstract
Violating the slow-roll regime during the final stages of inflation can significantly enhance curvature perturbations, a scenario often invoked in models producing primordial black holes and small-scale scalar induced gravitational waves. When perturbations are enhanced, one approaches the regime in which tree-level computations are insufficient, and nonlinear corrections may become relevant. In this work, we conduct lattice simulations of ultra-slow-roll (USR) dynamics to investigate the significance of nonlinear effects, both in terms of backreaction on the background and in the evolution of perturbations. Our systematic study of various USR potentials reveals that nonlinear corrections are significant when the tree-level curvature power spectrum peaks at $\mathcal{P}^{\rm max}_{\zeta}={\cal O}(10^{-3})-{\cal O}(10^{-2})$, with 5%--20% corrections. Larger enhancements yield even greater differences. We find a simple universal relation between simulation and tree-level quantities $\dot{\phi}=\dot{\phi}_{\rm tree}\left(1+\sqrt{\mathcal{P}^{\rm max}_{\zeta,\rm tree}}\right)$ at the end of the USR phase, which is valid in all cases we consider. Additionally, we explore how nonlinear interactions during the USR phase affect the clustering and non-Gaussianity of scalar fluctuations, crucial for understanding the phenomenological consequences of USR, such as scalar-induced gravitational waves and primordial black holes. Our findings demonstrate the necessity of going beyond leading order perturbation theory results, through higher-order or non-perturbative computations, to make robust predictions for inflation models exhibiting a USR phase.
@@ I Introduction
However, in certain models, the inflaton potential features a flat region or shallow minimum where the deceleration becomes dominant, and the field’s velocity can decrease exponentially, in what is called an “ultra-slow-roll” (USR) phase Ivanov et al. (1994); Kinney (1997); Inoue and Yokoyama (2002); Kinney (2005); Martin et al. (2013); Motohashi and Hu (2017). Perturbations generated during USR are far from scale-invariant, displaying a large amplification. These perturbations may lead to interesting phenomenological consequences, such as the emission of scalar-induced gravitational waves (SIGW) Tomita (1975); Matarrese et al. (1994); Acquaviva et al. (2003); Mollerach et al. (2004); Ananda et al. (2007); Baumann et al. (2007); Domènech (2021), and/or seed primordial black hole (PBH) formation Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974); Carr (1975); Chapline (1975). PBHs are a long-standing candidate for dark matter, potentially accounting for all dark matter or acting as seeds for supermassive black holes, see Refs. Carr et al. (2021); Green and Kavanagh (2021) for recent reviews. The blooming field of GW astronomy will provide unprecedented observations of small-scale GWs, being able to constrain the early universe (see Achúcarro et al. (2022); Auclair et al. (2023); Bagui et al. (2023) for recent reviews).
A very relevant question is whether the dynamics of USR are consistent with perturbativity (see Ref. Kristiano and Yokoyama (2024a) for a recent review). From a technical standpoint, the dimensionless power spectrum of curvature perturbations, $\mathcal{P}_{\zeta}(k)$, is usually calculated within linear, i.e. tree-level, perturbation theory. However, as linear perturbations can become large (often reaching $\sqrt{\mathcal{P}_{\zeta}(k)}\simeq 0.1$), curvature perturbations are expected to be sensitive to nonlinear interactions. These nonlinear effects have various consequences. First, they introduce nonzero higher-order cumulants in addition to the variance, with important implications for the PBH formation, see e.g. Franciolini et al. (2018); Taoso and Urbano (2021). Second, through loops, they modify the variance itself compared to the result obtained from the free theory. Lastly, they lead to a backreaction on the background dynamics. In this paper we investigate the relevance of these effects, considering motivated scenarios built with the reverse engineering approach proposed in Ref. Franciolini and Urbano (2022), by performing numerical lattice simulations able to capture the dynamics of the inflaton non-perturbatively.
@@ II.1 Inflationary background
The inflationary background can be described by modeling the evolution of the Hubble rate $H\equiv\dot{a}/a$, where $a$ is the scale factor during inflation. This is dictated by dynamical equations relating $H$ to the Hubble parameters, which are
@@
where $\dot{H}=dH/dt$ is the cosmic-time derivative of $H$ while $N$, defined as $dN=Hdt$, is the number of $e$-folds.111 Notice that a different definition of the Hubble parameter $\eta$ is sometimes used, and defined as $\epsilon_{2}\equiv\dot{\epsilon}/H\epsilon=-2(\eta-\epsilon)$.
@@
where the parameter $\delta N$ controls the width of the two transitions at $N_{\rm in}$ and $N_{\rm end}$. We assume that the second Hubble parameter $\eta$ is negligible in the initial slow-roll phase $\eta_{\rm I}\simeq 0$, consistent with the CMB requirement in the earlier stage of inflation. Once the parameters in Eq. (2) are fixed, one can solve for the evolution of $\epsilon$ and $H$ as a function of number of $e$-folds using (1). In all cases, we arbitrarily set $N_{\rm in}=0$. In Fig. 1, we show the evolution of $\epsilon$ and $\eta$ for the scenarios considered in this work, which can be divided into three cases, depending on the value of $\eta$ during USR: i) the approximate Wands duality (case I); ii) the repulsive (case II); iii) the attractive (case III). The choice of these names will become clear in the following.
@@
Figure 1: Evolution of $\eta(N)$ (top panel) and $\epsilon(N)$ (bottom panel) as a function of number of $e$-folds considered in this work. Solid, dashed and dotted lines indicate different cases, corresponding to the approximate Wands duality (WD), repulsive and attractive case, respectively. See the main text for more details. Different colors (from purple to yellow) correspond to different USR durations $N_{\rm end}-N_{\rm in}$, leading to a tree-level curvature power spectrum at its maximum ranging from ${\cal P}_{\zeta}^{\rm max}=10^{-4}$ to ${\cal P}^{\rm max}_{\zeta}=1$. In all cases, we fixed $N_{\rm in}=0$.
@@
In Tab. 1 we report parameters we adopt in the various cases. We arbitrarily fix $\delta N=0.5$, in order to have relatively smooth transitions in and out of the USR phase. Additionally, we choose values of $\eta_{\rm II}$ that are in all cases larger than $\simeq 3/2$, as required to enter in the USR regime. Finally, we set $\eta_{\rm III}=-0.5$, as a negative value is required in order to end inflation by raising $\epsilon$ again to reach ${\cal O}(1)$. [^2] For simplicity, we use the terminology “slow-roll” for the third phase, by contrast with the preceding USR phase, despite the relatively large value of $\eta_{\rm III}$. In case I, we approximately realise a scenario that respects the Wands duality (WD) Wands (1999), stating that the same tree-level spectrum of perturbations is obtained during phases characterised by a constant $\eta$ or $3-\eta$. In case I, this property is respected between phases II and III, and therefore it leads to a tree-level curvature power spectrum featuring a single smooth peak that can be fitted with a broken power-law. This happens because the spectrum of modes exiting the Hubble scale during USR follows $k^{2(3-\eta_{\rm II})}\sim 1/k$, while $k^{2\eta_{\rm III}}\sim 1/k$ in the subsequent phase Karam et al. (2023); Franciolini and Urbano (2022). This property can also be derived by inspecting the effective mass of scalar perturbations when $\epsilon$ is negligible. We will come back to this point in Sec. II.3.2. As we will see, the nearly constant effective mass term will also be accompanied by negligible higher-order derivatives of the potential, rendering the theory nearly free from self-interactions at these times. In cases II and III, $\eta_{\rm II}$ is taken to be smaller or larger than the WD value, leading to different phenomenology for the scalar perturbations.
@@ II.2 Inflaton potential
where in the second equation we consider the minus sign having in mind a large-field model in which the field value decreases as inflation proceeds. Combining $V(N)$ and $\phi(N)$, we reconstruct the profile $V(\phi)$ of the inflationary potential in field space. In all cases, we arbitrarily choose $\phi(N_{\rm CMB})=3.5$ and, by fixing the amplitude of the power spectrum at large scales to be compatible with CMB data, i.e. ${\cal P}_{\zeta}=2.1\cdot 10^{-9}$, we find $V(N_{\rm CMB})=3.2\cdot 10^{-9}$. The initial condition satisfying the SR attractor at $N_{\rm CMB}$ require $\phi(N_{\rm CMB})=3.5$ and ${\rm d}\phi(N_{\rm CMB})/{\rm d}N=-0.11$. Eq. (3a) shows the convenience of modeling the inflationary dynamics directly at the level of $\eta$ instead of $V(\phi)$. This is because the Hubble parameters enters at the exponent of the definition of $V(N)$, and thus allow for a much finer control on power spectral features when performing the reverse engineering procedure.
In Fig. 2, we show the reconstructed potential for the various scenarios considered in this work, zooming in the shallower region of the potential that controls the USR dynamics. We also show the first and second derivatives of the inflaton potential, as they will be relevant for the subsequent discussion. It is interesting to notice that, in cases II and III, sharp changes of the second derivative of the potential are obtained near the end of the feature, corresponding to the expected end of the USR phase. This will lead to large higher-order derivatives of the potential, crucially controlling the coupling of the theory. Notice that the potentials of case II are monotonic. Cases I and III present instead a local minimum followed by a local maximum (in the direction of decreasing $\phi$). However, the minimum is not deep enough to lead to a trapping of the inflaton Atal et al. (2019, 2020); Inomata et al. (2022); Escrivà et al. (2023); Caravano et al. (2024) (contrary to what happens in other contexts, see e.g. Caravano et al. (2024) in a model with oscillations in the potential Inomata et al. (2023)), with the exception of the most extreme situation of case III with $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=1$, as we will discuss below.
@@ II.3 Tree-level curvature power spectrum
Once the background evolution is defined, we can calculate the spectrum of gauge-invariant comoving curvature perturbations generated during inflation. As long as the slow-roll approximation holds, this spectrum is given by ${\cal P}_{\zeta}(k)={H^{2}}/{8\pi^{2}\epsilon}$ where the Hubble parameters are evaluated at the Hubble crossing of mode $k$. This hints toward the general expectation of exponentially enhanced perturbations during the USR phase, where $\epsilon$ decays exponentially.
@@
We solve the SM equation with sub-Hubble Bunch-Davies initial conditions for modes deep within the Hubble sphere, i.e. when $N$ is smaller than the Hubble crossing epoch $N_{k}$ such that $k=a(N_{k})H(N_{k})$. This is implemented on the SM variable as
@@
To illustrate the typical dynamics, Fig. 3 shows the evolution of curvature perturbations as a function of the number of $e$-folds for Case I with $\Delta N=2$ (corresponding to a very good approximation with $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-2}$). In the left panel, we show modes crossing the Hubble sphere before, during and right after the USR phase. While each mode experiences a growth due to the decaying mode changing behavior during USR, they all freeze out soon after USR ends and the subsequent SR evolution begins. The lattice simulations track the evolution of the system until the relevant modes have frozen. In the right panel, we indicate where these modes sit in the curvature power spectrum at the end of inflation. We normalize the x-axis with respect to $k_{\rm ref}$, which is the mode crossing the Hubble scale at $N_{\rm in}$, i.e. $k_{\rm ref}=a(N_{\rm in})H(N_{\rm in})$.
@@ II.3.1 On the irrelevance of gravitational backreaction
We can inspect the relevance of the gravitational backreaction by focusing on the SM equation (6). By neglecting gravitational backreaction, i.e. neglecting the lapse and shift perturbations in the spatially flat gauge, one obtains a different mass term, namely Eq. (7) subtracting $4\epsilon\eta-6\epsilon-2\epsilon^{2}$. In the situations of interest we have $\epsilon\ll 1$, which shows that gravitational backreaction can be safely neglected as long as $|3-\eta|\gg\epsilon$. This simply corresponds to the decoupling limit (see e.g. Creminelli et al. (2024)) in the effective field theory of inflation Cheung et al. (2008). This condition is satisfied in our setups and hence our lattice simulations, which do not take into account metric fluctuations, should agree with tree-level results if standard perturbation theory holds. [^3] Note that, if necessary, metric perturbation can be perturbatively taken into account in lattice simulations, as recently done in Caravano and Peloso (2024). We have confirmed this explicitly by checking the relative difference between the effective mass terms computed in the two cases. Due to the exponential decay of $\epsilon$ during the USR phase $\epsilon\sim\exp\left(-2\eta_{\rm II}N\right)$, the backreaction induced terms quickly become subdominant, while $m_{\rm eff}^{2}/H^{2}$ remains ${\cal O}(1)$.